OpenStudy (anonymous):

(q^(-2)-r^(-2))/q^(-1)-r^(-1))

OpenStudy (anonymous):

I know the answer is q+r/qr but I have now idea how they got the answer

OpenStudy (anonymous):

here we have to use two laws of exponents first is \[a ^{m}/a ^{n} = a ^{m-n}\] and \[a ^{-m} = 1/a ^{m}\]

OpenStudy (anonymous):

\[\frac{q^{-2}-r ^{-2}}{q ^{-1}-r ^{-1}}\]right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thats right @ cali

OpenStudy (anonymous):

\[q^{-2-(-1)} +r ^{-2-(-1)}\]

OpenStudy (anonymous):

\[q ^{-2+1}+r^{-2+1}\]

OpenStudy (anonymous):

\[q^{-1}+r^{-1}\]

OpenStudy (anonymous):

\[\frac{1}{q+r}\]

OpenStudy (anonymous):

so 1 1 r^2 - q^2 q^(-2) - r^(-2) = ------ - ------- = ---------------- q^2 r^2 q^2 * r^2 1 1 r - q and q^(-1) - r^(-1) = ------- - ------ = ----------- q r qr

OpenStudy (anonymous):

how did the denominator become a multiplication problem?

OpenStudy (anonymous):

Now putting these together we get q^(-2) - r^(-2) r^2 - q^2 q * r (r-q) (r+q) ( q*r) ---------------- = -------------- x ---------- = ---------------- q^(-1) - r^(-1) q^2 * r^2 r - q (q*r)(q*r) (r-q) cancelling out common factors we get q^(-2) - r^(-2) (r+q) r + q ---------------- = ----- = ----------- q^(-1) - r^(-1) (q*r) qr

OpenStudy (anonymous):

what do you mean chris by saying "denominator became a multiplication problem"??

OpenStudy (anonymous):

it was a subtraction problem, I'm sorry, you are righ on the money I'd give you 2 gold stars if I could and become a fan...I just don't understand

OpenStudy (anonymous):

ok i'll break it down further...

OpenStudy (anonymous):

i understand the cancelling at the end but how you change the q^-1-r^-1 into q^1 * r^1 i don't understand

OpenStudy (anonymous):

see numerator of the given problem is q^(-2) - r^(-2) q^(-2) = 1/q^2 and r^(-2) = 1/r^2 as per the law i gave so it becomes 1 1 q^(-2) - r^(-2) = ------ - ------- q^2 r^2 now this is of the ty[pe fraction - fraction and so we have to take LCD here since we do not have numbers, we take LCD as product of the terms in the denominator to get q^2*r^2 so then we get r^2 - q^2 q^(-2) - r^(-2) = ---------- q^2 * r^2 We do similarly for q^(-1) - r^(-1) do you understand now??????

OpenStudy (anonymous):

hold on let me read... lol your being a big help thanks

OpenStudy (anonymous):

ok, however if u hv a SKYPE account, v cud hv voice chat n it will b easier to explain

OpenStudy (anonymous):

yea, i'll check into it, but I don't have it yet but if I see you on again I'll definitely hit you up

OpenStudy (anonymous):

oooohhh that's the common denominator

OpenStudy (anonymous):

that's where I've been fkn up

OpenStudy (anonymous):

YEAH !!!!!!!

OpenStudy (anonymous):

yea, I totally understand now dude, wow thanks

OpenStudy (anonymous):

GOOD now give me the second medal..... :))

OpenStudy (anonymous):

wish I could

OpenStudy (anonymous):

here i got one for ya

OpenStudy (anonymous):

check the board lol

OpenStudy (anonymous):

where ??????